OpenStudy (anonymous):

4/9=x/45 how would you solve this proportion

pooja195 (pooja195):

\[\huge~\rm~\frac{ 4 }{ 9 }=\frac{ x }{45 }\] cross multiplcation

pooja195 (pooja195):

\[\huge~\rm~~4 \times 45=? \]\[\huge~\rm~9 \times x=?\]